A coffee maker is rated at 1170 W, a toaster at 1050 W, and a waffle maker at 1370 W. The three appliances are connected in parallel to a common 120 V household circuit.

(a) What is the current in each appliance when operating independently?
coffee maker A?
toaster A?
waffle maker A?

(b) What total current is delivered to the appliances when all are operating simultaneously?
A?

(c) Is a 15-A circuit breaker sufficient in this situation? Explain.
Answer #1

a) Power = voltage times current, hence I = P/V

- Coffee maker I = 1170W/120V = 9.75 A
- Toaster    I = 1050W/120V = 8.75 A
- Waffle maker I = 1370W/120V = 11.42 A

b) First you will need all the resistors, in order to create just on equivalent resistor and hence finding the total current.

Remember that R = P/i2

Rcoffee = 1170/(9.75)2= 12.31Ω

Rtoaster = 1050/(8.75)2=13.71Ω

Rwafflem = 1370/(11.42)2=10.51Ω

NOW, using because they are all parallel we can add them in the following way

Req = 1/[(1/R1)+(1/R2)+(1/R3)]

and hence Req = 4.01 Ω

NOw that we know the Req we can find the total current using V=RI

I=V/R = 120/4.01 = 29.93 A

c) A 15 A breaker is not sufficient because when all the appliances are on, the total current surpasses the 15 A.